OpenStudy (anonymous):

A 47-foot-long piece of pipe is to be cut into three pieces. The second piece is three times as long as the first piece, and the third piece is two feet more than five times the length of the first piece. What is the length of the longest piece?

OpenStudy (anonymous):

47 = x + 3x + (5x + 2) solve for x and then get (5x+2), the longest piece

OpenStudy (whpalmer4):

DWC is setting up the problem using \(x\) as the length of the first piece, as the lengths of the other pieces are defined in terms of the first piece. This is the most straightforward way to set up this problem, but you do have to remember that the answer to the problem is NOT the value of \(x\)!

OpenStudy (whpalmer4):

a seat-of-the-pants method of looking at this might be "all of the pieces are multiples of something, except the last one which is a multiple plus 2 feet. take that 2 feet off the total and we have 45 feet of pipe and 1 + 3 + 5 = 9 parts of the pipe, with the long piece being 5 parts (plus the 2 feet we set aside). 45/9 = 5 feet for each part, so the long piece is 5 parts of 5 feet each plus the 2 feet we set aside."

OpenStudy (anonymous):

A. 5 B.15 C.27 D.32 are the whats i have to pick from

OpenStudy (whpalmer4):

yes, the correct answer is one of those.